OpenStudy (anonymous):

Bill earns $12 per hour and works at most 40 hours per week. Identify the independent and dependent quantity in the situation, and find reasonable domain and range values.

OpenStudy (anonymous):

sorry id the anwser D:

OpenStudy (anonymous):

$12*(40) = ?

OpenStudy (anonymous):

480

OpenStudy (anonymous):

money earned = m(x) m(x) = (12)(x) x is the number of hours. x is the domain x must be greater than 0, but less than or equal to 40 hours per week m(x) is the range. it goes from 0 to (12*40)

OpenStudy (anonymous):

thank you for giving me a medal!! :D